# Sequential Model for Predicting Patient Adherence in Subcutaneous Immunotherapy for Allergic Rhinitis

### v2

@@ 4.2.3 LSTM
Long short-term memory (LSTM) is a classical sequential model (Hochreiter and Schmidhuber, 1997). The primary objective of this study is to forecast $y_{t}$ utilizing historical data, formulated as $y_{t}=f(x_{1:t},y_{1:t-1},s)$. To align this approach with the SLAC model for score prediction, an additional term $x_{t+1}$ is also predicted,
$(x_{t+1},y_{t})=f(x_{1:t},y_{1:t-1},s)$ (15)
where $f$ is a function represented by LSTM. The loss consists of cross entropy for adherence classification and Normalized Mean Squared Error Loss (NMSE) for score prediction.
In our scenarios, SLAC stands out due to its inherent flexibility over traditional sequential models like LSTM. This flexibility is primarily observed in its predictive capabilities. SLAC can predict $y_{t}$ and use this prediction to influence the subsequent $x_{t+1}$. In contrast, LSTM only predicts a pair of $y_{t}$ and $x_{t+1}$ simultaneously, implying that we cannot use $y_{t}$ to alter $x_{t+1}$. Although it is possible to modify the LSTM model to predict a pair of $y_{t}$ and $x_{t}$, this approach encounters a similar issue for $y_{t}$: it cannot predict $y_{t}$ using the information from $x_{t}$.
@@ 5 Results
The observed total non-adherence rate at the end of three years was 35.4% and the median of the SCIT duration was 18 months in the study. The rate of dropout in the third year(43.0%) was highest in comparison to the end of the first year $(26.5\,\%)$ and the second year $(30.0\,\%)$. The reason for the withdrawal from the patients included the concern of COVID-19, especially at the beginning of 2020 accounting for a $25\,\%$ portion of the non-adherent patients in the first year. The most influential reason for the withdrawal was unreached expectations for clinical improvement (43.9%). Medical issues including pregnancy status during the treatment period and other physical disorders were collected from patients leading to the withdrawal of SCIT $(4.5\,\%)$. The significantly improved symptoms contributed to the reason for dropout, especially after two years SCIT treatments $(18.2\,\%)$. The recorded cases from side effects accounted for $4.5\,\%$, comprising the local and systematic adverse reactions (see Table 2).
We have a total of 205 samples, which we have randomly divided into a test dataset comprising $20\,\%$, i.e., 41 samples. For our analysis, we employ a five-fold cross-validation approach. Additionally, we apply zero-mean and unit standard deviation (STD) normalization to the variables $x$ and $s$.
The Root Mean Square Error (RMSE) metric is used to evaluate the precision of our medical score predictions. Furthermore, to assess the adherence predictions, we utilize a comprehensive set of metrics including accuracy, precision, recall, and the F1 score, each offering a unique perspective on the performance of our predictive models. Most of the figures in this study are presented using boxplots.
The uncertainties for both models are calculated using five-fold cross-validation. In addition, as SLAC is a probabilistic model, we also perform 100 samples from the latent space to compute its uncertainty.
@@ 5.1 One-Step Prediction
SLAC $\mathbf{1.00\pm 0.00}$ $\mathbf{0.75\pm 0.03}$ $0.74\pm 0.03$ $0.71\pm 0.03$ $0.44\pm 0.05$
@@
SLAC $1.00\pm 0.01$ $\mathbf{0.87\pm 0.06}$ $\mathbf{0.90\pm 0.03}$ $\mathbf{0.86\pm 0.04}$ $\mathbf{0.70\pm 0.10}$
@@
In this experiment, our focus is on predicting the immediate next step. Within the SLAC, the prediction of $y_{t}$ is based on the sequence $x_{1:t}$ and actions $a_{1:t-1}$. Additionally, we forecast the subsequent state $x_{t+1}$ using the sequence $x_{1:t}$ along with actions $a_{1:t}$. In contrast, for LSTM, the predictions for both $y_{t}$ and the next state $x_{t+1}$ are derived from $x_{1:t}$ and $y_{1:t-1}$.
@@
As illustrated in Fig.3, SLAC surpasses LSTM in performance beginning at time step two. The figure indicates that with an increased amount of historical data (additional time steps), SLAC achieves greater RMSE. Both SLAC and LSTM demonstrate considerably better over random prediction methods. Further insights are provided in Fig. 5 and Fig. 6, which provides detailed representations of each feature. In the prediction of specific local symptoms score, SLAC performs an improvement in error after step two with all parameters compared to LSTM. The results from RMSE in nasal and ocular symptoms display relatively high values compared to those for lower respiratory tract symptoms. This can be attributed to the fact that the majority of patients in the cohort predominantly exhibited nasal and ocular symptoms, which presented a wide range of scores.
@@
Fig. 4 demonstrates that from steps two to four, accuracy in adherence predictions improves with the inclusion of additional information. The first step exhibits a notable bias, as it only includes data from adherent patients, as detailed in Sec. 4.1.2. Nonetheless, both models adeptly manage this bias and achieve high-accuracy predictions. Prediction for the sixth step is not conducted due to the cessation of treatment by the hospital. In the fifth step, there is a decline in accuracy, likely due to the extended time interval of 12 months. Table 3 illustrates details of the classification for one-step prediction. Initially, both models exhibit perfect performance in Accuracy, Precision, and Recall at the first time step, but diverge in subsequent steps. In terms of Accuracy, LSTM generally outperforms SLAC, particularly evident at time steps three, four, and five. For Precision, LSTM again shows superior performance in the later time steps, except at time step two where SLAC marginally leads. However, in the Recall metric, SLAC surpasses LSTM from time step two onwards, indicating its strength in correctly identifying positive cases. The F1 score, which balances precision and recall, shows LSTM generally ahead, except at time step two where SLAC has a slight edge. This metric indicates LSTM’s balanced capability in both precision and recall, especially in the later time steps. Overall, while both models start equally strong, LSTM demonstrates greater consistency and effectiveness across most metrics and time steps. SLAC, while lagging slightly behind in accuracy and precision, shows its robustness in recall, especially in the middle to later time steps.
@@ 5.4 Interpretability
5.5 Architecture and computation
The SLAC model’s architecture featured 32 hidden dimensions each for variables $z_{1}$ and $z_{2}$. Its encoder and decoder were symmetrically structured, each comprising five layers with 128 units. The primary activation function was LeakyReLU, set with a negative slope coefficient of 0.2. Both the encoder and decoder’s mean output layers were linear, while the STD layer utilized a Softplus activation. For binary classification tasks, a Sigmoid activation was used for output.
The LSTM architecture included a hidden dimension size of 128, with two LSTM layers. The output activation function for score prediction was linear, and like the SLAC model, a Sigmoid function was employed for binary classification outputs.
Both models shared the same optimization settings. They utilized the RAdam (Liu et al., 2019) optimizer with a learning rate of $0.001$. The batch size was set at 64, and a gradient clipping value of 0.8 was applied to ensure training stability. To prevent overfitting and enhance model generalization, a dropout rate of 0.05 was introduced. Additionally, both models incorporated Mixup as a data augmentation during training.
6 Discussion
@@
The reported adherence rates of SCIT ranged from around $23\,\%$ to $90\,\%$, due to the non-uniform follow-up duration (2-4 years) (Passalacqua et al., 2013; Lee et al., 2019; Lemberg et al., 2017; Yang et al., 2018). Poor adherence in the three to five-year time span of AIT is an obstacle to reaching allergen tolerance and symptom remission. Recently proposed an ‘adherence and persistence in AIT (APAIT)’ checklist that assists researchers in assessing adherence or persistence to AIT in retrospective studies (Pfaar et al., 2023). The present study is the first research about the application of machine learning models in the adherence prediction of SCIT in AR patients. From our study, the accomplishment rate of the three-year treatment cycle was relatively low ($35.4\,\%$), while the rate of dropout after two years accounts for half ($42.8\,\%$) in the whole non-adherence cohort. Several researchers focus on these variables impacting the compliance of medical behaviors to enhance the intervention approach to reduce the withdrawal caused by disease-unrelated reasons. Even though the Covid-19 pandemic impacted the compliance of the majority of the patients in the three-year cycle, the first year since the pandemic’s outbreak appears to fundamentally built a barrier to patients, similar to the finding from Liu et al. (2021) that $11\,\%$ dropouts in the two years SCIT was observed caused by Covid-19. We excluded the patients who dropped out in the dose buildup phase within four months to minimize the dose-origin impact. Due to the uncovered cost from the public health care system and commercial insurance, financial burden accounts for a non-negligible factor in influencing the decision-making of patients. A similar finding from Lourenço et al. (2020). indicated that economic reasons contributed to the most frequent cause of SCIT cessation.
In comparison to the logistic regression analyses in the clinical field to identify independent predictors, the model from our study performed convincing interpretability. Previous research primarily concentrated on non-sequential prediction methods for adherence (Mousavi et al., 2022; Wang et al., 2020; Warren et al., 2022; Ruff et al., 2019). This approach presents a significant limitation in treatment processes, particularly for immunotherapy that often spans extended periods, such as three years. These non-sequential methods tend to predict only the overall outcome, overlooking the intricacies of intermediate time steps. To facilitate earlier intervention, a sequential model capable of making predictions at any given time step would be markedly more beneficial. While some subsequent studies have introduced sequential models (Hsu et al., 2022; Singh et al., 2022; Schleicher et al., 2023), their scope was restricted to predicting adherence alone. Our study enhances this approach by incorporating a state-action model, which can predict both adherence and score/state. This advancement allows for more precise and detailed analysis of patient cases by medical professionals.
@@
Our study demonstrates notable findings in the domain of patient adherence prediction in subcutaneous immunotherapy. The comparison between the SLAC model and LSTM model reveals the distinct strengths and limitations of each approach. Notably, the SLAC exhibits greater flexibility, and it outperforms the LSTM in score prediction. This advantage likely stems from its ability to efficiently learn and generalize in complex environments, a trait that is crucial in medical data analysis. Conversely, the LSTM model shows better performance in predicting adherence, indicating its potential utility in scenarios where accurate forecasting of patient compliance is critical. Both models demonstrate the capability to handle longer sequences, extending beyond one-step prediction. This ability is crucial in medical settings where long-term patient monitoring and prediction are essential for effective treatment planning.
@@
Overall, the study underscores the importance of selecting the appropriate model based on the specific requirements of the task, whether it be flexibility, precision in score prediction, or adherence prediction. The findings contribute to the growing field of machine learning applications in healthcare, particularly in enhancing patient-centered treatment strategies through accurate and personalized predictions. Future research could focus on evaluating the SLAC model’s performance in simulating various actions, further enriching its applicability in clinical settings.
7 Conclusion
In summary, our study showcases the flexibility and score prediction superiority of the Sequential Latent Actor-Critic (SLAC) model, while highlighting the superior adherence prediction capabilities of the Long Short-Term Memory (LSTM) model. Both models demonstrate proficiency in managing extended sequence predictions, extending beyond single-step forecasts. This study emphasizes the utility and significance of selecting the appropriate model based on specific task requirements, be it for precision in score prediction or adherence forecasting. These findings make a substantial contribution to the application of machine learning in healthcare, particularly in improving patient-centered treatment strategies through accurate and tailored predictions.
@@ 7 Conclusion
The dataset for this study can be found in the Subcutaneous-Immunotherapy-Dataset repository.

### v3

@@ 3.2.3 LSTM
As an alternative, Long short-term memory (LSTM) is a classical sequential neural-network model (Hochreiter and Schmidhuber, 1997). Given the historical data, in our implementation, an LSTM predicts the score $x_{t+1}$ and adherence $y_{t}$ in parallel. The LSTM’s autoregressive feature allows us to iteratively input its current predictions to predict subsequent outcomes, covering prediction from $x_{t+1}$ to $x_{T}$ and $y_{t}$ to $y_{T-1}$. Implementation details and the reasons for the model choice can be found in Appendix B.
@@ 4 Results
The observed total non-adherence rate at the end of three years was 35.4% and the median of the SCIT duration was 18 months in the study. The rate of dropout in the third year (43.0%) was highest in comparison to the end of the first year (26.5%) and the second year (30.0%). The reason for the withdrawal from the patients included the concern of COVID-19, especially at the beginning of 2020 accounting for a 25% portion of the non-adherent patients in the first year. The most influential reason for the withdrawal was unreached expectations for clinical improvement (43.9%). Medical issues including pregnancy status during the treatment period and other physical disorders were collected from patients leading to the withdrawal of SCIT (4.5%). The significantly improved symptoms contributed to the reason for dropout, especially after two years SCIT treatments (18.2%). The recorded cases from side effects accounted for 4.5%, comprising the local and systematic adverse reactions (see Table 2).
We have a total of 205 samples, which we have randomly divided into a test dataset comprising 20%, i.e., 41 samples. For our analysis, we employ a five-fold cross-validation approach. Additionally, we apply zero-mean and unit standard deviation (STD) normalization to the variables $x$ and $s$.
The Root Mean Square Error (RMSE) metric is used to evaluate the precision of our medical score predictions. Furthermore, to assess the adherence predictions, we use a comprehensive set of metrics including accuracy, precision, recall, and the F1 score, each offering a unique perspective on the performance of our predictive models. Most of the figures in this study are presented using boxplots.
In all results in this study, the uncertainties for both models are calculated using five-fold cross-validation. In addition, as SLAC is a probabilistic model, we also perform 100 samples from the latent space to compute its uncertainty.
@@ 4.1 One-Step Prediction
SLVM $\mathbf{1.00\pm 0.00}$ $\mathbf{0.75\pm 0.03}$ $0.74\pm 0.03$ $0.71\pm 0.03$ $0.44\pm 0.05$
@@
SLVM $1.00\pm 0.01$ $\mathbf{0.87\pm 0.06}$ $\mathbf{0.90\pm 0.03}$ $\mathbf{0.86\pm 0.04}$ $\mathbf{0.70\pm 0.10}$
@@
In this experiment, our focus is on predicting the immediate next step. Within the SLVM, the prediction of $y_{t}$ is based on the sequence $x_{1:t}$ and actions $a_{1:t-1}$. Additionally, we forecast the subsequent state $x_{t+1}$ using the sequence $x_{1:t}$ along with actions $a_{1:t}$. In contrast, for LSTM, the predictions for both $y_{t}$ and the next state $x_{t+1}$ are derived from $x_{1:t}$ and $y_{1:t-1}$.
@@
As illustrated in Fig. 3, SLVM surpasses LSTM in performance beginning at time step two. The figure indicates that with an increased amount of historical data (additional time steps), SLVM achieves greater RMSE. Both SLAC and LSTM demonstrate considerably better over random prediction methods. Further insights are provided in Fig. 5 and Fig. 6, which provides detailed representations of each feature. In the prediction of specific local symptoms score, SLVM performs an improvement in error after step two with all parameters compared to LSTM. The results from RMSE in nasal and ocular symptoms display relatively high values compared to those for lower respiratory tract symptoms. This can be attributed to the fact that the majority of patients in the cohort predominantly exhibited nasal and ocular symptoms, which presented a wide range of scores.
@@
Fig. 4 demonstrates that from steps two to four, accuracy in adherence predictions improves with the inclusion of additional information. The first step shows a notable bias, as it only includes data from adherent patients, as detailed in Sec. 3.1.2. Nonetheless, both models adeptly manage this bias and achieve high-accuracy predictions. Prediction for the sixth step is not conducted due to the cessation of treatment by the hospital. In the fifth step, there is a decline in accuracy, likely due to the extended time interval of 12 months. In future research, it would be worthwhile to explore whether adopting a consistent interval for data collection could enhance the outcomes of longitudinal prediction. Table 3 illustrates details of the classification for one-step prediction. Initially, both models exhibit perfect performance in Accuracy, Precision, and Recall at the first time step, but diverge in subsequent steps. In terms of Accuracy, LSTM generally outperforms SLVM, particularly evident at time steps three, four, and five. For Precision, LSTM again shows superior performance in the later time steps, except at time step two where SLVM marginally leads. However, in the Recall metric, SLVM surpasses LSTM from time step two onwards, indicating its strength in correctly identifying positive cases. The F1 score, which balances precision and recall, shows LSTM generally ahead, except at time step two where SLAC has a slight edge. This metric indicates LSTM’s balanced capability in both precision and recall, especially in the later time steps. Overall, while both models start equally strong, LSTM demonstrates greater consistency and effectiveness across most metrics and time steps. SLVM, while lagging slightly behind in accuracy and precision, shows its robustness in recall, especially in the middle to later time steps.
@@ 4.4 Interpretability
5 Discussion and Conclusion
@@
The reported adherence rates of SCIT ranged from around 23% to 90%, due to the non-uniform follow-up duration (2–4 years) (Passalacqua et al., 2013; Lee et al., 2019; Lemberg et al., 2017; Yang et al., 2018). Poor adherence in the three to five-year time span of AIT is an obstacle to reaching allergen tolerance and symptom remission. Recently “adherence and persistence in AIT (APAIT)” checklist was proposed to assists researchers in assessing adherence or persistence to AIT treatment(Pfaar et al., 2023). The present study is the first research regarding the application of machine learning models in the adherence prediction of SCIT in AR patients. From our study, the accomplishment rate of the three-year treatment cycle was relatively low (35.4%), while the dropout rate after two years accounts for half (42.8%) in the whole non-adherence cohort. Several researchers focus on these variables impacting adherence to medical behaviors to enhance the intervention approach to reduce the withdrawal caused by disease-unrelated reasons. Even though the COVID-19 pandemic affected the adherence of the majority of the patients in the three-year cycle, the first year since the pandemic’s outbreak appears to have fundamentally built a barrier to patients, similar to the finding from Liu et al. (2021) that 11% dropouts in the two years SCIT was observed caused by COVID-19. We excluded the patients who dropped out in the dose buildup phase within four months to minimize the dose-origin impact. Due to the uncovered cost from the public health care system and commercial insurance, financial burden accounts for a non-negligible factor in influencing the patient’s decision-making. A similar finding from Lourenço et al. (2020) indicated that economic reasons contributed to the most frequent cause of SCIT cessation.
Recently, as the presented studies focusing on the medication adherence prediction of non-communicable diseases such as diabetes, hypertension, cancer, and chronic respiratory diseases regarding machine learning models were introduced into the application, the systematic monitoring of patients’ adherence behaviors remarkably re-tailored the disease management and enhanced medical decision-making (Kanyongo and Ezugwu, 2023) Due to the multidimensional variables in the prediction of adherence collected for analysis, machine learning models exhibit advantages in automatic feature selection, interaction effects, scalability, robustness, and so on compared to traditional regression analysis.
Mousavi et al. (2022) demonstrated the effectiveness of a hybrid model that combines neural networks and genetic algorithms for predicting diet adherence. Wang et al. (2020) explored another hybrid model that integrates neural networks and support vector machines to predict nonadherence in Crohn’s Disease patients by streamlining the intervention process in medicine-taking. Both methods have shown excellent performance. The deployment of machine-learning algorithms in the prediction of adherence in cardiovascular disease including random forests, support vector machines, and neural networks showed the accuracy ranged from 0.53 to 0.97 (Zakeri et al., 2022; Mirzadeh et al., 2022). The ensemble learning model in the prediction of adherence from the patients who conducted self-administer injections proposed by Gu et al. (2021) achieved a good performance and generalization properties based on the fusion of multiple heterogeneous classifiers. In the field of allergen immunotherapy, Yao et al. (2023) introduced a machine-learning model with an improved DFSSA algorithm to predict the therapeutic efficacy of AIT for asthma using clinical characteristics and serum allergen detection metrics. However, these non-sequential methods generally predict only the final outcome, neglecting the complexities of intermediate stages. A sequential model that can make predictions at any specific time step would significantly enhance the ability for early intervention. Hsu et al. (2022) investigated the advantages of incorporating patient history into the prediction of medication adherence. They assessed the performance of temporal neural network models, particularly LSTM and simple recurrent neural networks, and compared these with non-temporal neural networks, ridge classifiers, and logistic regression. To optimize the efficacy of cognitive training for older adults, Singh et al. (2022) employed multivariate time series analysis and developed personalized models for each patient to capture their unique adherence patterns. However, the sequential data of patients is often characterized by fluctuating adherence and high dropout rates, resulting in uneven, unaligned, and missing values in the time series data. To address this challenge, Schleicher et al. (2023) applied change point detection to identify phases with varying dropout rates, presented methods for handling uneven and misaligned time series, and used time series classification to predict the user’s phase. These models, however, overlook the significance of score prediction in SCIT treatment. Our study advances this methodology by integrating a state-action model capable of predicting both adherence and score/state. This enhancement facilitates a more accurate and comprehensive whole-process management of AR patients in SCIT treatment.
The proposed prediction models can help clinicians dynamically measure the effectiveness of adherence interventions including more frequent reminders or engagement strategies, such that healthcare teams can focus on these individuals and proactively provide them with additional support. Moreover, in the surveillance of local symptom scores and rescue medicine up-take, the models offer a clinical evidence-based approach to precisely predict the risk of non-adherence in patient-centered care precisely. Especially for the potential risk of withdrawal caused by medical issues or related side effects, our models suggest multidimensional observational parameters for timely offering medical intervention and increasing patient engagement by participating in shared decision-making. The implementation of integration of the models with existing healthcare workflows is challenging, while the application of telesystem and online consultation would improve the work efficacy in immunotherapy centers and facilitate patient’s self-management.
@@
Our study demonstrates notable findings in the domain of patient adherence prediction in subcutaneous immunotherapy. The comparison between the SLAC model and LSTM model reveals the distinct strengths and limitations of each approach. Notably, SLAC exhibits greater flexibility, and it outperforms LSTM in score prediction. This advantage likely stems from its ability to efficiently learn and generalize in complex environments. Conversely, the LSTM model shows better performance in predicting adherence, indicating its potential usage in scenarios. Both models demonstrate the capability to handle longer sequences, extending beyond one-step prediction. This ability is crucial in medical settings where long-term patient monitoring and prediction are essential for effective treatment planning.
@@
Overall, the study underscores the importance of selecting the appropriate model based on the specific requirements of the task, whether it be flexibility, precision in score prediction, or adherence prediction. The findings contribute to the growing field of machine learning applications in healthcare, particularly in enhancing patient-centered treatment strategies through accurate and personalized predictions. Future research could focus on evaluating the SLAC model’s performance in simulating various actions, further enriching its applicability in clinical settings. Additionally, the generalization to other diseases or the application of our models would be an interesting direction for future research.
@@
The dataset for this study can be found in a GitHub repository https://github.com/leexxe/Subcutaneous-Immunotherapy-Dataset.
Appendix A Latent sequential variable model
[…]
figures/lvm
Figure 10: Schematic of the LSVM part of SLAC. Solid and dashed lines denote the generative and inference model pathways, respectively. The gray circles represent observed data, and the white circles denote latent variables. The figure is adapted from (Lee et al., 2020).
The sequential latent variable model (SLVM) of the SLAC consists of an inference model and a generative model (see Fig. 10). The inference model in a sequential latent-variable model typically aims to approximate the posterior distribution of the latent variables given the observed data. It tries to infer the hidden states $z$ based on the observed inputs $x$ and initial states $s$. The inference models the probability distributions of the latent variables $z^{1}$ and $z^{2}$ at different time steps. $q_{\phi}$ denotes the variational distribution parameterized by $\phi$,
$z^{1}_{1}$ $\sim q_{\phi}(z_{1}^{1}\mid x_{1},s)$ (1)
$z^{2}_{1}$ $\sim p_{\phi}(z^{2}_{1}\mid z_{1}^{1})$ (2)
$z^{1}_{t+1}$ $\sim q_{\phi}(z^{1}_{t+1}\mid x_{t+1},z^{2}_{t},a_{t})$ (3)
$z^{2}_{t+1}$ $\sim p_{\phi}(z^{2}_{t+1}\mid z^{1}_{t+1},z^{2}_{t},a_{t}).$ (4)
The generative model, on the other hand, describes how the observed data is generated from the latent variables. The generative model is the probability distribution of both the initial latent states and their transitions over time, as well as the likelihood of the observations given the latent states, with $p_{\phi}$ indicating the parameterized generative distribution.
$z^{1}_{1}$ $\sim p(z_{1}^{1})$ (5)
$z^{2}_{1}$ $\sim p_{\phi}(z^{2}_{1}\mid z_{1}^{1})$ (6)
$z^{1}_{t+1}$ $\sim p_{\phi}(z^{1}_{t+1}\mid z^{2}_{t},a_{t})$ (7)
$z^{2}_{t+1}$ $\sim p_{\phi}(z^{2}_{t+1}\mid z^{1}_{t+1},z^{2}_{t},a_{t})$ (8)
$x_{t}$ $\sim p_{\phi}(x_{t}\mid z_{t}^{1},z^{2}_{t})$ (9)
$y_{t}$ $\sim p_{\phi}(y_{t}\mid z_{t}^{1},z^{2}_{t}).$ (10)
We have the evidence lower bound (ELBO):
$\log p_{\phi}(x_{1:t+1}|a_{1:t})\geq$ $\bigg{[}\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\bigg{[}\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\sum^{T-1}_{t=0}\Big{(}\log p_{\phi}(x_{t+1}\mid z_{t+1})$ (11)
$-D_{KL}\bigl{(}q_{\phi}(z_{t+1}\mid x_{t+1},z_{t},a_{t})\;\|\;p_{\phi}(z_{t+1}\mid z_{t},a_{t})\bigr{)}\Big{)}\bigg{]}\bigg{]}.$
For ease of notation, we have $q(z_{1}\mid x_{1},z_{0},a_{0}):=q(z_{1}\mid x_{1},s)$ and $p(z_{1}\mid z_{0},a_{0}):=p(z_{1})$. The ELBO provides a lower bound to the log-likelihood of the observed data, which is computationally intractable to compute directly. It is composed of two terms: the expected log-likelihood of the observed data given the latent variables, and the Kullback-Leibler (KL) divergence between the variational distribution and the prior distribution of the latent variables. Minimizing the KL divergence can be interpreted as enforcing the variational distribution to be as close as possible to the prior, while maximizing the expected log-likelihood ensures that the model accurately captures the distribution of the observed data. To predict the adherence, we have $\log p_{\phi}(y_{t+1}|z_{t+1})$ as a regularisor in the loss function.
$\min_{\phi}$ $\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\left[\sum^{T-1}_{t=0}\left[D_{KL}\left(q_{\phi}(z_{t+1}\mid x_{t+1},z_{t},a_{t})\;\|\;p_{\phi}(z_{t+1}\mid z_{t},a_{t})\right)\right]\right]$ (12)
s.t. $\quad\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\left[\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\left[\sum^{T-1}_{t=0}-\log p_{\phi}(x_{t+1}\mid z_{t+1})\right]\right]\leq\xi_{\mathrm{score}}$ (13)
$\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1},y_{1:T-1})\sim D}\left[\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\left[\sum^{T-2}_{t=0}-\log p_{\phi}(y_{t+1}\mid z_{t+1})\right]\right]\leq\xi_{\mathrm{adherence}}$ (14)
where $\xi$ is a baseline error, in Eq. (13) we have regression with Gaussian distribution, and in Eq. (14) we use cross-binary entropy loss for classification. To solve the optimization problem, we incorporate the constraints into the objective function using Lagrange multipliers $\lambda$. We apply methods from (Chen et al., 2022) to adapt $\lambda$. This allows the model to balance the importance of the constraints relative to the divergence terms, which can help in avoiding common pitfalls in training such as suboptimal local minima and posterior collapse.
Appendix B LSTM
The primary objective of this study is to forecast $y_{t}$ from historical data, formulated as $y_{t}=f(x_{1:t},y_{1:t-1},s)$. To align this approach with the SLVM of SLAC for score prediction, an additional term $x_{t+1}$ is also predicted,
$(x_{t+1},y_{t})=f(x_{1:t},y_{1:t-1},s)$ (15)
where $f$ is a function represented by an LSTM. The loss consists of the cross entropy for adherence classification and the Normalized Mean Squared Error Loss (NMSE) for score prediction.
In our scenarios, SLVM stands out due to its inherent flexibility over traditional sequential models like LSTM. This flexibility is primarily observed in its predictive capabilities. SLVM can predict $y_{t}$ and use this prediction to influence the subsequent $x_{t+1}$. In contrast, LSTM only predicts a pair of $y_{t}$ and $x_{t+1}$ simultaneously, implying that we cannot use $y_{t}$ to alter $x_{t+1}$. Although it is possible to modify the LSTM model to predict a pair of $y_{t}$ and $x_{t}$, this approach encounters a similar issue for $y_{t}$: it cannot predict $y_{t}$ using the information from $x_{t}$.
Appendix C Architecture and computation
The SLVM model’s architecture featured 32 hidden dimensions each for variables $z_{1}$ and $z_{2}$. Its encoder and decoder were symmetrically structured, each comprising five layers with 128 units. The primary activation function was LeakyReLU, set with a negative slope coefficient of 0.2. Both the encoder and decoder’s mean output layers were linear, while the STD layer utilized a Softplus activation. For binary classification tasks, a Sigmoid activation was used for output.
The LSTM architecture included a hidden dimension size of 128, with two LSTM layers. The output activation function for score prediction was linear, and as in the SLVM model, a Sigmoid function was used for binary classification outputs.
Both models shared the same optimization settings. They used the RAdam (Liu et al., 2019) optimizer with a learning rate of 0.001. The batch size was set at 64, and a gradient clipping value of 0.8 was applied to ensure training stability. To prevent overfitting and enhance model generalization, a dropout rate of 0.05 was introduced. Additionally, both models incorporated Mixup as a data augmentation during training.